OpenStudy (anonymous):

solve the equations by using the square root property or completing the square: 1.) y^2=8 2.)(9s+8)^2=9 show steps / solution

OpenStudy (anonymous):

Use the square root property.

OpenStudy (anonymous):

You forgot the +/-. \[y^2 = 8 \implies y = \pm \sqrt{8} = \pm \sqrt{2^2 \cdot 2} = \pm 2\sqrt{2}\]

OpenStudy (anonymous):

1.) y^2=8 \[y=\sqrt{8}=2\sqrt{2}\]

OpenStudy (anonymous):

i have the answers i need the solutions

OpenStudy (anonymous):

use caculater if you want exactly answer

myininaya (myininaya):

approximate answer?

OpenStudy (anonymous):

ey nancy i think its + or - because you can square it

OpenStudy (anonymous):

Yeah, the above is the exact answer, the calculator will give you the approximate one.

OpenStudy (anonymous):

2.)(9s+8)^2=9\[9s+8=+-\sqrt{9}\]

OpenStudy (anonymous):

\[s=(-8\pm \sqrt{8})/9\]

OpenStudy (anonymous):

they got the answers

OpenStudy (anonymous):

\[s=\frac{-8+\sqrt{8}}{9}\]or\[s=\frac{-8-\sqrt{8}}{9}\]caculator